OpenStudy (anonymous):

Which of the following is the ratio of circumference to diameter? radius center circle pi

OpenStudy (ipwnbunnies):

π Is the ratio of the circumference to the diameter of a circle: \[Circumference = \pi d\] \[\pi = \frac{circumference}{diameter}\]